OpenStudy (anonymous):

find a power series representation of (6t)/(5-t^(5))

OpenStudy (anonymous):

can anyone solve this problem?

OpenStudy (anonymous):

I might be able to, no promises. Give me a sec

OpenStudy (anonymous):

do you have a solution? from the back of the book or something

OpenStudy (anonymous):

okay I got the answer and I think its right, you there?

OpenStudy (anonymous):

yeah sorry. I dont have an answer though...

OpenStudy (anonymous):

okay I got...

OpenStudy (anonymous):

\[6t * \sum_{n=0}^{\infty} (t ^{5}/5)^{n}\]

OpenStudy (anonymous):

is that what the question is asking? And see if you can see how I got that, if you can't I'll try and explain

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think it should be 6t/5 *∑n=(t5/5)^n...

OpenStudy (anonymous):

he's right I have that written down on paper but didn't write it when coppying to computer

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

great - thank you

OpenStudy (anonymous):

*Virtual high five*

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Inik take a look at the question that I posted...

OpenStudy (anonymous):

estimate the value...?

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

medal pwease yeah inik that one

OpenStudy (anonymous):

I have another question coming up just like this one

OpenStudy (anonymous):

Cool. Are you gunna post a new one or keep on this